The region in the first quadrant, which is bounded by the curve


x² = 4y, the line x = 4, is revolved about the line x = 4. Locate the


centroid of the resulting solid of revolution.

1
Expert's answer
2022-06-09T14:55:34-0400
"x=4: 4^2=4y=>y=4"

"x^2=4y, x\ge0=>: x=2\sqrt{y}, y\ge 0"


"V=\pi\displaystyle\int_{0}^{4}(4^2-(2\sqrt{y})^2) dy"

"=\pi[16y-2y^2]\begin{matrix}
 4 \\
 0
\end{matrix}=32\pi"

In this case, the centroid lies on the axis "x=4."

"V_y=\pi\displaystyle\int_{0}^{4}y(2\sqrt{y}) dy"

"=\pi[\dfrac{4y^{5/2}}{5}]\begin{matrix}
 4 \\
 0
\end{matrix}=\dfrac{128\pi}{5}"

"\bar{y}=\dfrac{\dfrac{128\pi}{5}}{32\pi}=0.8"